OpenStudy (anonymous):

what is the nth term to 18 11 4 -3 -10

OpenStudy (akashdeepdeb):

This is an Arithmetic sequence. 18,11,4,-3,-10.......................n Now do you see the pattern. It REDUCES BY 7 So common difference = -7 First term = 18 n = ? An = a + (n-1)d = 18 + (n-1)*-7 = 18 - 7n + 7 = -7n + 25 This will give you the value of ANY term 'n' in the sequence! Got it? :)